OpenStudy (anonymous):

For 1 month, two doctors in a medical office counted the number of patients they saw each day. The mean and MAD for each doctor were calculated and recorded in a table. What conclusion can be drawn, based upon the value of the means-to-MAD ratio? Doctor Mean MAD Hartmann 15.3 0.9 Kim 18.9 1.1 A. Dr. Hartmann saw more patients per day than Dr. Kim did. B. Dr. Kim saw more patients per day than Dr. Hartmann did. C. Dr. Kim and Dr. Hartmann saw about the same number of patients per day. D. No conclusions can be drawn.

OpenStudy (anonymous):

doctor MEAN MAD H 15.3 0.9 K 18.9 1.1

OpenStudy (anonymous):

@Michele_Laino @parakeet @Bookworm14 @maryfitzgerald14 @IrishBoy123 @jim_thompson5910 @amistre64 @EclipsedStar @wio @RockerChic8 @jpgonzales55 @Loser66 @Music_me

OpenStudy (anonymous):

i am not sure about this plz help!

OpenStudy (anonymous):

i will give medal and fan :)

OpenStudy (anonymous):

@perl @tkhunny

OpenStudy (anonymous):

i am think'in either b or c

OpenStudy (anonymous):

not d for sure

OpenStudy (anonymous):

no, i am think b now

OpenStudy (michele_laino):

I think that we have to use the ration: mean over MAD

OpenStudy (michele_laino):

oops..ratio*

OpenStudy (anonymous):

haha, kk

OpenStudy (michele_laino):

for example: case Dr. Hartmann mean/MAD= 15.3/0.9= 17 case Dr. Kim mean/MAD= 17.2 approximately

OpenStudy (tkhunny):

As a good general rule, don;'t just post and tag. Show your work. Anything you can give us will help.

OpenStudy (anonymous):

can u help me with this please?!?!

OpenStudy (anonymous):

@Bookworm14

OpenStudy (michele_laino):

I think that we have to use the ration: mean over MAD

OpenStudy (anonymous):

@Bookworm14 , do u have the time right now, i am desperate...

OpenStudy (anonymous):

thanks for com'in :)

OpenStudy (bookworm14):

What is the MAD? lol im in highschool and ive never seen that before? lol and your welcome, sorry i couldn't get on earlier, ive got ALOT of work to do :(

OpenStudy (anonymous):

it is okay, so in other, words, r u joking or do u really don't know?!!

OpenStudy (bookworm14):

I am not 100% sure, but I know how to do the problem, i think lol so mean "to" mad means divide your mean (15.3) by your MAD (0.9) etc etc so Dr. Hartman seen 15.3/0.9 = 17 patients Dr. Kim seen 18.9 / 1.1= 17.18 repeated patients. so they both aw about 17 patients, or about the same amount each so what should the answer be? :)

OpenStudy (anonymous):

if u do not know this one, then can u help me with something else?

OpenStudy (anonymous):

oh, oky u know, lol

OpenStudy (anonymous):

okay, i get the concept thx! do u mind helping me with maybe one more? just one

OpenStudy (bookworm14):

the answe is C, and sorry, but I really don't have the time :( please don;t be mad, im just slammed with work right now :(